OpenStudy (r_josh10):

I will tag you for help. will fan and medal. @Nurali @TheEdwardsFamily @danish071996 and anyone else can help

OpenStudy (r_josh10):

Here is the question.

OpenStudy (anonymous):

do cross multiply

OpenStudy (anonymous):

then solve the equation and try to find out x/y

OpenStudy (r_josh10):

umm im confused. sorry

OpenStudy (nurali):

3/5=x/x+2y cross multiply 5x=3(x+2y) 5x=3x+6y 5x-3x=6y 2x=6y x/y=6/2 x/y=?

OpenStudy (anonymous):

nurali will help u

OpenStudy (r_josh10):

3? @Nurali

OpenStudy (nurali):

Correct